OpenStudy (anonymous):

Suppose JK bisects angle LJM. The measure of angle LJK = (-10x + 3) and the measure of KJM = (-x + 21)º. Which equation represents the best way to find the measure of angle LJM? (Hint: Sketch the situation first.)

OpenStudy (anonymous):

 LJK = KJM -10x +3 = -x + 21 -10x + x = 21 - 3 -9x = 18 x = -2 KJM = -x + 21 = 2 + 21 = 23 LJM = LJK + KJM LJM = 23 + 23 LJM = 46º